Sam and Jane Hill, both age 35, are married filing a joint return. Jane is employed full time and Sam is a part owner in several local businesses. They have contacted you inquiring about the Section 199A qualified business income (QBI) deduction. They have provided information for their Year 1 business income in the exhibit above. Sam and Jane do not elect to aggregate any of the qualifying businesses. Their only other income in Year 1 is Jane's salary of $250,000. They will not be itemizing deductions in Year 1 and will use the married filing jointly (MFJ) standard deduction of $24,400. For purposes of this simulation, ignore any deduction for one-half of self-employment tax.

QBI DEDUCTION FOR EACH BUSINESS

TAXABLE INCOME

TAXALBE INCOME BEFORE QBI DEDUCTION

QBI DEDUCTION FOR EACH BUSINESS

APLHA PARTNERS

BRAVO COMPANY INC

CHARLIE COMPANY LLC

TOTAL SECTION 199A QBI DEDUCTION

TAXABLE INCOME AFTER QBI DEDUCTION


Alpha
Partners

Bravo
Co. Inc.

Charlie Co.
LLC

Total

Ownership

Sam
—25%

Sam
—25%

Sam
—25%

General Partner

Shareholder

Member-Manager

Entity form

Partnership

S Corp.

Limited Liability Co.

Type of business

Bookstore

Gift Shop

Coffee Shop

Share of ordinary business income (loss)

($20,000)

$50,000

$150,000

$180,000

Salary or guaranteed payment

$15,000

$30,000

$45,000

Withdrawals/distributions

$5,000

$10,000

$15,000

Other Information Provided by Business:

Total W-2 wages paid by the business

$40,000

$30,000

$220,000

$290,000

Total unadjusted basis immediately after acquisition (UBIA) of qualified property

199A QBI Taxable Income Limitations

Filing Status

Taxable Income

Single and all other

$160,700–210,700

Married filing jointly

$321,400–$421,400

Using the exhibits, calculate the amount of Sam and Jane's qualified business income (QBI) deduction for each business and the total Section 199A QBI deduction and enter the amounts in the table below. Use whole numbers, and if the amount of the deduction is zero, enter a zero (0).
Answer #1

199A = 20% of Qualified Business Income

Salary or Guaranteed Payment is not qualified for 199A.

Sam total income = 180,000 + 45,000 = 225,000 + 250,000(Jane) = 475,000 Which is more than 421,400. for MFJ

Charlie Co.LLC deduction will be lessor of below:

i. 20% of QIB (150,000) = 30,000

II. 50% of Wages paid to Employee (220,000) = 110,000

So 30,000 will be 199A deduction for Charlie Co.LLC

Bravo Co

i. 20% of QIB (50,000) = 10,000

II. 50% of Wages paid to Employee (30,000) = 15,000

So 10,000 will be 199A deduction for Charlie Co.LLC.

Alpha Parnters amount will be Zero.

Taxable income after QIB

Total Income - 475,000 - 10,000 - 30,000 = 435,000

TAXALBE INCOME BEFORE QBI DEDUCTION

475,000

QBI DEDUCTION FOR EACH BUSINESS

APLHA PARTNERS

NIL

BRAVO COMPANY INC

10,000

CHARLIE COMPANY LLC

30,000

TOTAL SECTION 199A QBI DEDUCTION

40,000

40,000

TAXABLE INCOME AFTER QBI DEDUCTION

435,000